Calculate the volume of carbon(IV)oxide evolved at s t.p by dissolving 20grams of pure marble,CaCO3 in dilute trioxonitrate(v)acid
1
Expert's answer
2020-03-31T12:30:16-0400

CaCO3 + 2 HNO3 = Ca(NO3)2 + CO2 + H2O

There were 20 g/100 g/mole = 0.2 mole of CaCO3 and equimolar amount of CO2 evolved.

CO2 gas volume evolved at STP is 0.2 mole*22.4 L/mole = 4.48 L